OpenStudy (piasolywoda):

where is (cosx - cos2x)/sinx divergent?

OpenStudy (solomonzelman):

I feel the questions lacks some content.

OpenStudy (terryhall1999):

Z 5 2 dx √ 3x − 6 = 6 (convergent) (1) Z ∞ 0 dx x 2 + 4 = π 4 (convergent) (2) Z 1 0 ln(x) dx = −1 (convergent) (3) Z ∞ 0 e x e 2x + 1 dx = π 4 (convergent, u = e x ) (4) Z ∞ −∞ x x 4 + 1 dx = 0 (convergent) (5) Z ∞ 0 xex dx (divergent)